# Imitation from Observations with Trajectory-Level Generative Embeddings

### v1

@@ 4 Methodology
We aim to develop a novel offline imitation learning from observations framework that leverages trajectory-level generative planners as encoders. Such planners enable the algorithm to reason over extended horizons and incorporate rich temporal structure that is difficult to capture using stepwise imitation approaches. To achieve this, we propose Trajectory-level Generative Embeddings (TGE), a method that constructs reward signals from the latent embeddings of a generative planner trained to perform density estimation over both the suboptimal dataset $\mathcal{D}_{\mu}$. These trajectory-level latent embeddings provide a discriminative representation capable of distinguishing expert-like and suboptimal trajectories, thereby serving as an informative reward proxy for offline policy optimization on $\mathcal{D}_{\mu}$. Finally, we train an offline RL algorithm on the suboptimal dataset using the TGE-derived reward function. The overall procedure is summarized in Algorithm 1, and an overview of the pipeline is illustrated in Figure 1.
4.1 Generative Planner Learns Separable Embeddings
Recent work has shown that diffusion models can be viewed through the lens of denoising autoencoders across multiple noise scales, and that intermediate diffusion features can serve as useful representations for downstream tasks fuest2024diffusion. In particular, yang2023diffusion propose to extract and distill multi-timestep diffusion representations (RepFusion) and demonstrate strong performance on recognition benchmarks, suggesting that diffusion training induces semantically meaningful feature hierarchies.
Motivated by this perspective, we treat the encoder of the trajectory diffusion planner Diffuser janner2022planning as an unsupervised representation learner and use its trajectory-level latent embedding for imitation. Specifically, as outlined in Line 1 of Algorithm 1, we train the diffusion noise model on the suboptimal dataset $\mathcal{D}_{\mu}$ using the objective
$\mathcal{L}(\theta)=\mathbb{E}_{\tau_{0}\sim\mathcal{D}_{\mu},\,k,\,\epsilon}\bigl[\lVert\epsilon-\epsilon_{\theta}(\tau_{k},k)\rVert_{2}^{2}\bigr],$ (1)
where $k\sim\mathrm{Unif}[K]$ denotes the diffusion timestep, $\epsilon\sim\mathcal{N}(0,I)$ is the Gaussian noise target, and $\tau_{k}$ is the noised trajectory obtained by adding noise to the clean trajectory $\tau_{0}$ with length $H$ at diffusion timestep $k$. Let $\tau\in\mathcal{T}$ denote a state-action trajectory of length $H$, where the trajectory space is defined as $\mathcal{T}=\mathbb{R}^{H\times(|\mathcal{S}|+|\mathcal{A}|)}$. The noise model $\epsilon_{\theta}$ is parameterized using a U-Net architecture (ronneberger2015u), which we conceptually decompose into a temporal encoder $\phi:\mathcal{T}\times\mathbb{R}\to\mathcal{Z}$ and a decoder $\psi:\mathcal{Z}\times\mathbb{R}\to\mathcal{T}$, such that $\epsilon_{\theta}(\tau_{k},k)=(\phi_{\theta}~\circ~\psi_{\theta})(\tau_{k},k)$. Minimizing the denoising score-matching objective in Equation (1) encourages the latent representation $\mathcal{Z}$ to preserve information about the underlying clean trajectory distribution, as it must support reconstruction across a range of Gaussian noise levels. Consequently, $\mathcal{Z}$ captures temporally coherent features that are stable under perturbations, yielding an embedding space that is more semantically discriminative for downstream similarity-based rewards.
Given a clean input trajectory $\tau_{0}\sim\mathcal{D}_{\mu}$, we extract a latent embedding using only the encoder. Figure 2 illustrates that, even without any explicit discriminative supervision between expert-like and suboptimal trajectories in $\mathcal{D}_{\mu}$, the encoder-induced embedding $z=\phi_{\theta}(\tau_{0},0)$ naturally separates expert-like samples from suboptimal ones in the mixed dataset. Given this desirable separation property of the embeddings, it is natural to leverage them for reward construction by directly measuring an appropriate distance between embeddings.
4.2 Reward Estimation on Trajectory-level Embeddings
We now describe the procedure for generating synthetic rewards based on the trajectory-level embeddings induced by the encoder of the generative planner. Given the expert dataset $\mathcal{D}_{E}$, we first extract a set of expert embeddings $\mathcal{Z}_{E}=\{z_{E}=\phi_{\theta}(\tau)\mid\tau\in\mathcal{D}_{E}\}$ from the state-only expert trajectories (Line 2 in Algorithm 1). For each state $s_{t}$ in the suboptimal dataset $\mathcal{D}_{\mu}$, we consider the trajectory segment that follows it, $\tau_{\mu}=s_{t:t+H}$, and encode it into a latent vector $z_{t}=\phi_{\theta}(\tau_{\mu})$. In the expert embedding space $\mathcal{Z}_{E}$, we compute the $L_{2}$-norms between $z_{t}$ and all expert embeddings $z_{E}\in\mathcal{Z}_{E}$, and then identify the top-$m$ nearest neighbors, denoted by $\mathcal{N}_{m}(\mathcal{Z}_{E})$. This nearest-neighbor retrieval serves as a proxy for density estimation; we provide the theoretical derivation linking this metric to particle-based entropy maximization liu2021behavior in Section 4.3.
To estimate how likely a suboptimal trajectory segment $\tau_{\mu}$ resembles an expert trajectory, we apply a Parzen–Rosenblatt estimator on the distances between $z_{t}$ and its $m$ nearest expert neighbors:
where $\sigma$ is a kernel temperature. Note that we normalize all latent embeddings by projecting them onto the unit sphere (dividing each vector by its $L_{2}$-norm), ensuring that all embeddings have equal length and allowing for consistent distance computations under the kernel. We adopt a Laplacian kernel $f(d)=\exp(-d)$ rather than the standard Gaussian kernel, as we observe empirically that Gaussian kernels $\exp{(-d^{2})}$ decay too aggressively in high-dimensional latent spaces, resulting in overly sparse rewards on suboptimal data. In contrast, the heavier tail of the Laplacian kernel provides denser gradients that more effectively guide the agent from suboptimal regions toward the expert subspace.
4.3 Downstream Policy Learning
With these synthetic rewards, the augmented dataset $\mathcal{D}^{\prime}_{\mu}=\{(s_{t},a_{t},r_{\text{TGE}}(s_{t}))\}$ can be used to train any offline RL method for policy learning. In our experiments, we adopt two state-of-the-art offline RL algorithms, Implicit Q-Learning (IQL) (kostrikov2021offline) and Revisited Behavior Regularized Actor-Critic (ReBRAC) (tarasov2023revisiting), and demonstrate that both methods effectively learn high-quality policies when equipped with our synthetic reward signal.
Algorithm 1 Trajectory-Level Generative Embeddings
0: Offline suboptimal dataset $\mathcal{D}_{\mu}$, expert dataset $\mathcal{D}_{E}$, horizon $H$, number of candidates $m$, temperature $\sigma$, and a kernel function $f(\cdot)$.
1: Train generative planner on $\mathcal{D}_{\mu}$ and extract the encoder $\phi_{\theta}$.
2: Compute embeddings $\mathcal{Z}_{\mu}=\phi_{\theta}(\mathcal{D}_{\mu})$ and $\mathcal{Z}_{E}=\phi_{\theta}(\mathcal{D}_{E})$ on trajectory segments length $H$.
4:for each $z_{t}\in\mathcal{Z}_{\mu}$ do
5:  Compute distances $d(z_{t},z_{E})=\|z_{t}-z_{E}\|_{2}$ for all $z_{E}\in\mathcal{Z}_{E}$.
6:  Construct the top-$m$ nearest neighbor $\mathcal{N}_{m}(\mathcal{Z}_{E})$.
7:  $r_{\text{TGE}}(s_{t})\leftarrow\frac{1}{m}\sum_{j\in\mathcal{N}_{m}(\mathcal{Z}_{E})}f(-d(z_{t},z_{E})/\sigma)$.
8:end for
9: Run any offline RL algorithm on the suboptimal dataset augmented with the synthetic rewards, i.e., $\mathcal{D}^{\prime}_{\mu}=\{(s_{t},a_{t},r_{\text{TGE}}(s_{t}))\}$.
4.4 Reward Interpretation and Learning Signals
An Entropy Perspective on the Reward Formulation
@@ 4.4 Reward Interpretation and Learning Signals
It is important to note that conducting policy learning by maximizing the reward defined in Eq. 2 is equivalent to minimizing an approximation of the particle-based cross-entropy between the latent expert distribution and the behavioral distribution. Specifically, given $n$ data points $\{z_{i}\}_{i=1}^{n}$ sampled from a latent state distribution induced by a policy $\pi$, we leverage non-parametric entropy estimation techniques singh2003nearest; liu2021behavior to express the particle-based cross-entropy with respect to the expert distribution as:
$H(\rho_{\pi},\rho_{E})=-\frac{1}{n}\sum_{i=1}^{n}\log\left(\frac{m}{nv_{i}^{m}}\right)+b(m)\propto\sum_{i=1}^{n}\log v_{i}^{m},$
@@
$H(\rho_{\pi},\rho_{E})\propto\sum_{i=1}^{n}\log\|z_{i}-z_{E}^{m}\|^{d_{z}}.$
This entropy estimator indicates that minimizing cross-entropy requires minimizing the distances between the agent’s states and their expert neighbors. Given that the reward function we applied $f(d)=e^{-d}\approx 1-\log(1+d)$, optimizing the policy to maximize the reward in Eq. 2 is similar with minimizing the cross-entropy between the latent distributions of the expert and behavioral policies.
Analysis of Learning Signals
Building on the entropy-based interpretation of our reward, we analyze the structural differences between our generative metric learning approach and the prevailing distribution-matching paradigm.
@@
Dual optimization methods fundamentally rely on estimating density ratios, $w(s)\approx d^{\pi_{E}}(s)/d^{\mathcal{D}_{\mu}}(s)$, to reweight offline transitions. This formulation typically assumes a nontrivial degree of support overlap ma2022versatile, such that expert transitions are reasonably represented within the offline dataset. In more realistic settings where the offline data is largely disjoint from the expert distribution, this ratio can become ill-conditioned. For suboptimal states outside the expert support, $d^{\pi_{E}}(s)\to 0$, causing importance weights to vanish. As a result, supervision concentrates on a small subset of in-support states, while much of the offline data receives nearly indistinguishable near-zero weights, limiting the ability to provide graded feedback on how different suboptimal behaviors compare or how they should be improved.
@@
In contrast, TGE reduces its reliance on strict support overlap by constructing a distance-shaped reward in a learned latent space. Rather than assigning near-uniformly negligible supervision to out-of-distribution states, the exponential kernel $r(s)\propto\exp(-d(z_{s},z_{E})/\tau)$ induces a smoothly decaying, graded signal as a function of distance. This allows many suboptimal trajectories to remain comparable through their relative proximity to expert behavior, providing a usable preference signal that prioritizes trajectories closer to the expert instead of restricting learning to a small overlapping subset.
@@
The effectiveness of this extrapolation depends on the semantic quality of the underlying metric. Standard distance metrics, such as $L_{2}$ distance on raw observations, often fail due to state aliasing, where states that are close under a raw metric may be functionally distinct (e.g., spatially proximate states separated by an obstacle in maze environments). By leveraging the encoder of a trajectory-level diffusion model, $\phi_{\theta}$, we exploit the model’s generative inductive bias, which is trained to capture joint temporal structure over trajectories. As a result, the induced latent space is organized according to long-horizon dynamics rather than superficial static features, making the resulting distances more reflective of functional similarity and better suited for robust imitation learning.

### v2

@@ 4 Methodology
We aim to develop a novel offline imitation learning from observations framework that leverages trajectory-level generative planners as encoders. Such planners enable the algorithm to reason over extended horizons and incorporate rich temporal structure that is difficult to capture using stepwise imitation approaches. To achieve this, we propose Trajectory-level Generative Embeddings (TGE), a method that constructs reward signals from the latent embeddings of a generative planner trained to perform density estimation over the suboptimal dataset $\mathcal{D}_{\mu}$. These trajectory-level latent embeddings provide a discriminative representation capable of distinguishing expert-like and suboptimal trajectories, thereby serving as an informative reward proxy for offline policy optimization on $\mathcal{D}_{\mu}$. Finally, we train an offline RL algorithm on the entire offline suboptimal dataset $\mathcal{D}_{\mu}$ using the TGE-derived reward function. The overall procedure is summarized in Algorithm 1, and an overview of the pipeline is illustrated in Figure 1.
4.1 An Entropy Perspective on Distribution Matching
@@
The fundamental goal of imitation learning is to recover a policy $\pi$ such that its induced latent state distribution $\rho_{\pi}$ matches the expert distribution $\rho_{E}$. We formalize this by seeking to minimize the particle-based cross-entropy $H(\rho_{\pi},\rho_{E})$. Given $n$ data points $\{z_{i}\}_{i=1}^{n}$ sampled from a latent state distribution induced by a policy $\pi$, we leverage non-parametric entropy estimation techniques Singh et al. (2003); Liu and Abbeel (2021) to express the particle-based cross-entropy with respect to the expert distribution as:
$H(\rho_{\pi},\rho_{E})=-\frac{1}{n}\sum_{i=1}^{n}\log\Big(\frac{m}{nv_{i}^{m}}\Big)+b(m)\propto\sum_{i=1}^{n}\frac{\log v_{i}^{m}}{n},$
@@
$H(\rho_{\pi},\rho_{E})\!\approx\!\frac{1}{n}\sum_{i=1}^{n}\log\!\Big(\!c{+}\frac{1}{m}\!\!\sum_{z_{E}^{m}\in\mathcal{N}_{m}(\mathcal{Z}_{E})}\!\!\!\!\!\!\!\!\|z_{i}-z_{E}^{m}\|^{d_{z}}\!\Big).$ (1)
This derivation reveals that minimizing the cross-entropy between distributions is equivalent to minimizing the log-distance between behavioral samples and their expert neighbors. To provide a smooth and robost learning signal, we generalize this estimator by averaging over the entire $m$-nearest neighborhood.
Following Liu and Abbeel (2021), we approximate the estimator $\log(c+\frac{1}{m}\sum\|z_{i}-z_{E}\|^{d_{z}})$ as $\log(1+\frac{1}{m}\sum\|z_{i}-z_{E}\|)$ for numerical stability and define the surrogate reward by
$\textstyle{r(z_{i})=\log\left(1+\frac{1}{m}\sum_{z_{E}^{m}\in\mathcal{N}_{m}(\mathcal{Z}_{E})}\|z_{i}-z_{E}^{m}\|\right),}$ (2)
plugging this into (1) yields that the cross-entropy between the occupancy measure $\rho_{\pi}$ and the expert behavior $\rho_{E}$ can be approximated by:
$H(\rho_{\pi},\rho_{E})$ $\approx\frac{1}{n}\sum_{i=1}^{n}\log\Big(1+\frac{1}{m}\!\sum_{z_{E}^{m}\in\mathcal{N}_{m}(\mathcal{Z}_{E})}\!\!\!\!\!\!\!\|z_{i}-z_{E}^{m}\|^{d_{z}}\Big)$
$\approx\tfrac{1}{n}{\sum_{i=1}^{n}}r(z_{i})\approx{\mathbb{E}}_{z_{i}\sim\rho_{\pi}}[r(z_{i})]$
$=\mathbb{E}_{\pi}\left[{\sum_{h=0}^{\infty}}\gamma^{h}r(z_{h})\right],$ (3)
where the second line is because $z_{i}$ is sampled from $\rho_{\pi}$ and the third line holds according to the definition of occupancy measure $\rho_{\pi}(\cdot)=\sum_{h=0}^{\infty}\gamma^{h}P_{h}(\cdot|\pi)$. The entropy estimation (3) suggests that optimizing the policy to maximize the cumulative surrogate reward defined in (2) is equivalent to minimizing the cross-entropy between the latent distributions of the expert and behavioral policies. Having established this theoretical framework, we describe our method in the following sections.
4.2 Generative Planner Learns Separable Embeddings
We treat the encoder of the trajectory diffusion planner Diffuser Janner et al. (2022) as an unsupervised representation learner and use its trajectory-level latent embedding for imitation. As discussed in Section 2.3, the denoising objective encourages the model to learn temporally coherent features that are robust to noise, making them ideal for imitation. Specifically, as outlined in Line 1 of Algorithm 1, we train the diffusion noise model on the suboptimal dataset $\mathcal{D}_{\mu}$ using the objective
$\mathcal{L}(\theta)=\mathbb{E}_{\tau_{0}\sim\mathcal{D}_{\mu},\,k,\,\boldsymbol{\epsilon}}\bigl[\lVert\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\tau_{k},k)\rVert_{2}^{2}\bigr],$ (4)
where $k\sim\mathrm{Unif}[K]$ denotes the diffusion timestep, $\epsilon\sim\mathcal{N}(0,I)$ is the Gaussian noise target, and $\tau_{k}$ is the noisy trajectory obtained by adding noise to the clean trajectory $\tau_{0}$ with length $H$ at diffusion timestep $k$. Let $\tau\in\mathcal{T}$ denote a trajectory segment of length $H$, where the trajectory space for training is defined as $\mathcal{T}=\mathbb{R}^{H\times(|\mathcal{S}|+|\mathcal{A}|)}$. To construct these segments from $\mathcal{D}_{\mu}$, we extract overlapping sequences of length $H$ using a sliding window. While the diffusion model is trained on these state–action trajectories, the encoder $\phi_{\theta}$ is designed to produce a state-only latent embedding. Specifically, during the embedding process, we mask the action dimensions of $\tau$, ensuring the representation space $\mathcal{Z}$ captures temporal state dynamics consistent with the observation-only expert data in $\mathcal{D}_{E}$. We achieve this by conceptually decomposing the U-Net noise model $\epsilon_{\theta}$ (Ronneberger et al., 2015) into a temporal encoder $\phi:\mathcal{T}\times\mathbb{R}\to\mathcal{Z}$ and a decoder $\psi:\mathcal{Z}\times\mathbb{R}\to\mathcal{T}$, such that $\epsilon_{\theta}(\tau_{k},k)=(\psi_{\theta}~\circ~\phi_{\theta})(\tau_{k},k)$. Minimizing the denoising score-matching objective in Equation (4) encourages the latent representation $\mathcal{Z}$ to preserve information about the underlying clean trajectory distribution, as it must support reconstruction across a range of Gaussian noise levels. Consequently, $\mathcal{Z}$ captures temporally coherent features that are stable under perturbations, yielding an embedding space that is more semantically discriminative for downstream similarity-based rewards.
Given a clean input trajectory $\tau_{0}\sim\mathcal{D}_{\mu}$, we extract a latent embedding using the encoder. Figure 2 illustrates that, even without any explicit discriminative supervision between expert-like and suboptimal trajectories in $\mathcal{D}_{\mu}$, the encoder-induced embedding $z=\phi_{\theta}(\tau_{0},0)$ naturally separates expert-like samples from suboptimal ones in the mixed dataset. Given this desirable separation property of the embeddings, it is natural to leverage them for reward construction by directly measuring an appropriate distance between embeddings as we discussed in Sec. 4.1.
4.3 Reward Estimation on Trajectory-level Embeddings
We now describe the procedure for generating synthetic rewards based on the trajectory-level embeddings induced by the encoder of the generative planner. Given the expert dataset $\mathcal{D}_{E}$, we extract expert embeddings $\mathcal{Z}_{E}$ by applying a sliding window with a stride of 1 to the state-only expert episode. Specifically, $\mathcal{Z}_{E}=\{\phi_{\theta}(\tau_{i:i+H})\mid\tau\in\mathcal{D}_{E},0\leq i\leq|\tau|-H\}$, where the action dimensions of the input to $\phi_{\theta}$ are treated as null for these observation-only segments (Line 2 in Algorithm 1). For each state $s_{t}$ in the suboptimal dataset $\mathcal{D}_{\mu}$, we consider the trajectory segment followed by, $\tau_{\mu}=s_{t:t+H}$, and encode it into a latent vector $z_{t}=\phi_{\theta}(\tau_{\mu})$. In the expert embedding space $\mathcal{Z}_{E}$, we compute the $\ell_{2}$-norms between $z_{t}$ and all expert embeddings $z_{E}\in\mathcal{Z}_{E}$, and then identify the top-$m$ nearest neighbors $\mathcal{N}_{m}(\mathcal{Z}_{E})$.
To construct an entropy-based surrogate reward signal from the suboptimal trajectory segment $\tau_{\mu}$, we compute the average logarithmic distance between the latent embedding $z_{t}$ and its $m$ nearest expert neighbors:
where $\sigma$ is a kernel temperature. Note that we normalize all latent embeddings by projecting them onto the unit sphere (dividing each vector by its $L_{2}$-norm), ensuring that all embeddings have equal length and allowing for consistent distance computations under the kernel. We adopt a logarithmic kernel $f(d)=-\log(1+d)$. This serves as a proxy for density estimation, consistent with the theoretical derivation linking this metric to particle-based entropy maximization Liu and Abbeel (2021) provided in Section 4.1. To justify our theoretical insights, we provide the ablation study in Appendix C which shows that our theoretically grounded selection provides the best performance.
4.4 Downstream Policy Learning
With these synthetic rewards, the augmented dataset $\mathcal{D}^{\prime}_{\mu}=\{(s_{t},a_{t},r_{\text{TGE}}(s_{t}))\}$ can be used to train any offline RL method for policy learning. In our experiments, we adopt two state-of-the-art offline RL algorithms, Implicit Q-Learning (IQL) (Kostrikov et al., 2021) and Revisited Behavior Regularized Actor-Critic (ReBRAC) (Tarasov et al., 2023), and demonstrate that both methods effectively learn high-quality policies when equipped with our synthetic reward signal. Notably, for ReBRAC, we introduce an adaptive behavior cloning mechanism that weights the imitation penalty by the TGE reward to prioritize expert-like transitions during training. Full implementation details are provided in Appendix A.2.
Algorithm 1 Trajectory-Level Generative Embeddings
0: Offline suboptimal dataset $\mathcal{D}_{\mu}$, expert dataset $\mathcal{D}_{E}$, horizon $H$, number of candidates $m$, temperature $\sigma$, and kernel function $f(\cdot)$.
1: Train generative planner on $\mathcal{D}_{\mu}$ and encoder $\phi_{\theta}$.
2: Compute embeddings $\mathcal{Z}_{\mu}=\phi_{\theta}(\mathcal{D}_{\mu})$ and $\mathcal{Z}_{E}=\phi_{\theta}(\mathcal{D}_{E})$ on trajectory segments of length $H$.
4:for each $z_{t}\in\mathcal{Z}_{\mu}$ do
5:  Compute $d(z_{t},z_{E})=\|z_{t}-z_{E}\|_{2}$ for all $z_{E}\in\mathcal{Z}_{E}$.
6:  Identify the top-$m$ nearest neighbors $\mathcal{N}_{m}(\mathcal{Z}_{E})$.
7:  $r_{\text{TGE}}(s_{t})\leftarrow\frac{1}{m}\sum_{z_{nb}\in\mathcal{N}_{m}(\mathcal{Z}_{E})}f\left(d(z_{t},z_{nb})/\sigma\right)$.
8:end for
9: Run any offline RL algorithm on the suboptimal dataset augmented with synthetic rewards: $\mathcal{D}^{\prime}_{\mu}=\{(s_{t},a_{t},r_{\text{TGE}}(s_{t}),s^{\prime}_{t})\}$.
4.5 Learning under Support Mismatch
@@
We analyze the structural differences between our proposed TGE approach and the prevailing occupancy matching methods. Standard dual optimization fundamentally rely on estimating density ratios, $w(s)\approx d^{\pi_{E}}(s)/d^{\mathcal{D}_{\mu}}(s)$, to reweight offline transitions. This formulation typically assumes a nontrivial degree of support overlap Ma et al. (2022), such that expert transitions are reasonably represented within the offline dataset. In more realistic settings where the offline data poorly covers the expert distribution, this ratio can become ill-conditioned. For suboptimal states outside the expert support, $d^{\pi_{E}}(s)\to 0$, causing importance weights to vanish. As a result, supervision concentrates on a small subset of in-support states, while much of the offline data receives nearly indistinguishable near-zero weights, limiting the ability to provide graded feedback on how different suboptimal behaviors compare or how they should be improved.
@@
In order to reduce the reliance on strict support overlap, we seek to construct a distance-shaped reward in the learned latent space of the generative planner under the entropy formulation mentioned above. Rather than assigning near-uniformly negligible supervision to out-of-distribution states, the logarithmic entropy-based reward $r(s)=-\log(1+\|z_{s}-z_{E}\|)$ induces a smoothly decaying, graded signal that persists even for states far from the expert support. This allows suboptimal trajectories to remain comparable through their relative proximity to expert behavior, providing a dense preference signal that prioritizes trajectories closer to the expert instead of restricting learning to a small overlapping subset.
@@
The effectiveness of this extrapolation depends on the semantic quality of the underlying metric. Standard distance metrics, such as $L_{2}$ distance on raw observations, often fail due to state aliasing, where states that are close under a raw metric may be functionally distinct (e.g., spatially proximate states separated by an obstacle in maze environments). By leveraging the encoder of a trajectory-level generative planner, we exploit the model’s generative inductive bias, which is trained to capture joint temporal structure over trajectories. As a result, the induced latent space is organized according to long-horizon dynamics rather than superficial single-step features, making the resulting distances more reflective of functional similarity and suited for robust imitation learning.